OpenStudy (anonymous):

This will probably seem rather simple, but I am horrible at math unless it's directly applied. I'm looking at f(x)=-7x-5 and I'm supposed to graph it.

OpenStudy (anonymous):

Plug in values for x. Ex, if x=-1 = -7(-1)-5=2 Coordinate would be (-1,2)... keep plugging in different values for x.

OpenStudy (anonymous):

Equations of the form y=mx+b where m and b are arbitrary constants are called linear functions and will output a line when graphed. An easy way to know how to graph it is that b will be the value of y where it intersects the y axis (since that's where x=0) and m will represent the slope (ratio of rise and run). In your case, the line will intersect the y-axis at y=-5, or (0,-5) and will have a slope of -7 meaning that for every unit right you go, you will go down seven units. I highly recommended web applications like wolframalpha.com to check your graphing.

OpenStudy (anonymous):

The equation of a line is y=mx+b, where m is the slope and b is the y intercept. So with a -7 slope, the line should be dramatically downwards. With a b of -5, the line crosses the y axis at (0,-5). To find other points, just plug in values of x and solve for y, or f(x).

OpenStudy (anonymous):

Thanks =P